FIND THE NO OF REAL SOLUTIONS OF
x^2-3|x|+2=0

Grade:12th pass

3 Answers

Bevkoof Singh
10 years ago

iit should be 4

Vikas TU
10 years ago

x^2-3|x|+2=0

Case 1: when x > 0

x2 - 3x + 2 = 0

 

x2  - 2x - x + 2 = 0

x(x-2) -1(x-2) = 0

(x-1) (x-2) = 0

x = 1 and x = 2

 

Case 2: when x < 0

x2 + 3x + 2 = 0

x2 + 2x + x + 2 = 0

x(x+2) + 1(x+2) = 0

(x+1) (x+2) = 0

x = -1 and x = -2

 

From both cases x belongs to {-1, -2, 1, 2}

Hence it has four no. of solutions.

 

PLZ  APPROVE!

Prajwal kr
10 years ago

Take two cases:

1.x>=0:

x2-3x+2

Soln: 2 solutions.(x=2,1)

2.x<0

x2+3x+2

Solution: 2 solutions(x=-2,-1)

 

Therefore 4 solutions.